Turbid water is muddy or cloudy water. Sunlight is necessary for most life forms; thus turbid water is considered a threat to wetland ecosystems. Passive filtration systems are commonly used to reduce turbidity in wetlands. Suspended solids are measured in mg/l. Is there a relation between input and output turbidity for a passive filtration system and, if so, is it statistically significant? At a wetlands environment in Illinois, the inlet and outlet turbidity of a passive filtration system have been measured. A random sample of measurements are shown below. (Reference: EPA Wetland Case Studies.)

Reading 1 2 3 4 5 6 7 8 9 10 11 12
Inlet (mg/l) 9.0 55.1 26.2 11.4 16.0 64.5 22.0 57.9 37.9 37.2 29.6 37.3
Outlet (mg/l) 2.7 15.2 3.4 2.6 3.8 18.8 3.5 17.3 15.8 9.8 14.0 13.8

Use a 1% level of significance to test the claim that there is a monotone relationship (either way) between the ranks of the inlet readings and outlet readings.

(a) Rank-order the inlet readings using 1 as the largest data value. Also rank-order the outlet readings using 1 as the largest data value. Then construct a table of ranks to be used for a Spearman rank correlation test.

Reading Inlet
Rank x
Oulet
Rank y
d = x - y d2
1
2
3
4
5
6
7
8
9
10
11
12












































Σd2 =

(c) Compute the sample test statistic. (Use 3 decimal places.)
Answer #1


(a) Outlet Reading Inlet X d = | | | 8| ਗਗ p| 107 11 | 1 Cd =18 (c)R. =-bgd nen 6(18) 12-12 =1 - 108 1716 =1-0.063 = 0.937t=?:Vn-2 (0.937) V12-2 V1-(0.937) _2.963 0.349 = 8.490 The value of the sample test statistic is t = 8.490.